OpenStudy (ammarah):

You are rolling a 20-sided die where the sides are numbered 1 through 20. Find the indicated odds. In favor of rolling a 10

OpenStudy (ammarah):

@tkhunny @surjithayer @wolf1728 @wio

OpenStudy (tkhunny):

1:19 -- Why is that a mystery?

OpenStudy (ammarah):

wait how is it 1:19?

OpenStudy (ammarah):

@tkhunny

OpenStudy (tkhunny):

The PROBABILITY of rolling a 10 is 1/20. There are 20 possibilities and you want only one of them. The ODDS of getting a 10 are 1:19. There are twenty possibilities and you want only one of them, and there are 19 things you don't want.

OpenStudy (ammarah):

i stil dont get it... i need another example how about in favor of rolling a number less than 6 do u just do 5/20?

OpenStudy (tkhunny):

The PROBABILITY of rolling less than 6 is 5/20 = 1/4. There are 20 possibilities and you want five of them. The ODDS of getting less than 6 are 5:15 = 1:3. There are twenty possibilities and you want five of them, and there are 15 things you don't want.

OpenStudy (tkhunny):

There is some oddity associated with the words FOR or AGAINST. The odds FOR getting less than 6 are 5:15. The odds AGAINST getting less than 6 are 15:5. Personally, I don't care for that distinction, but you do have to take care to see that you know what someone else is saying. This does contribute to communication.

OpenStudy (ammarah):

i dont get it what did u divide?-

OpenStudy (tkhunny):

I counted. It's just counting. It's also important to remember how to reduce fractions and ratios, but that is not a significant part of the result. 20:24 = 10:12 = 5:6 -- It's all the same. It's easier to understand what happened if you do NOT reduce. It only maters what your application needs.

OpenStudy (ammarah):

oh how about against rolling a 1,3,5

OpenStudy (ammarah):

would it be 1:20, 3:20, 1:4?

OpenStudy (ammarah):

I have anotyher one against rolling a number greater thatn 13 I did 13/20 amd o gpt 13:20

OpenStudy (ammarah):

In favor of rolling an even number less than 10 i got 2:5

OpenStudy (tkhunny):

The PROBABILITY of rolling 1, 3, or 5 is 3/20. There are 20 possibilities and you want three of them. The ODDS of getting 1, 3, or 5 are 3:17. There are twenty possibilities and you want three of them, and there are 17 things you don't want.

OpenStudy (tkhunny):

The PROBABILITY of rolling greater than 13 is 7/20. There are 20 possibilities and you want seven of them. The ODDS of getting greater than 13 are 7:13. There are twenty possibilities and you want seven of them, and there are 13 things you don't want.

OpenStudy (tkhunny):

The odds AGAINST rolling greater than 13 are 13:7.

OpenStudy (tkhunny):

The PROBABILITY of rolling even and less than 10 is 4/20. There are 20 possibilities and you want four of them. The ODDS of rolling even and less than 10 are 4:16. There are twenty possibilities and you want four of them, and there are 16 things you don't want.

OpenStudy (tkhunny):

You MUST see the relationship between probability and odds. They are NOT the same thing. The probability denominator was ALWAYS 20. Never fails. This is not the case for odds. For odds, the denominator is the stuff you don't want.

OpenStudy (ammarah):

thankyou that makes more sense!

OpenStudy (ammarah):

can u help me with some more

OpenStudy (ammarah):

im opening a new question

OpenStudy (tkhunny):

Not if I just have to keep typing the same words over and over. :-) Actually, it's past my bedtime. Post to the public boards. There are others who can respond.

OpenStudy (ammarah):

wait how aboutif its against rolling an odd number gretaer than 10

OpenStudy (ammarah):

i got 9:10

OpenStudy (tkhunny):

Now, that would be just typing the same words over and over. Think on it. Greater than 10 is 11-20. That's 10 numbers. Go!!

OpenStudy (tkhunny):

9+10 = 19. You're missing one.

OpenStudy (ammarah):

but its says agianst!

OpenStudy (tkhunny):

When they are the same, it maters not. 10:10 is the same as 10:10.

OpenStudy (ammarah):

so the answers 10:10

OpenStudy (tkhunny):

Is it FOR or AGAINST?

OpenStudy (ammarah):

against

OpenStudy (tkhunny):

Trick question. THEY are the SAME! It doesn't matter. Just seeing if you are paying attention. gtg

OpenStudy (ammarah):

so i was right?!!//

OpenStudy (tkhunny):

Don't act so surprise. Surely it happens from time to time. Good work.

OpenStudy (ammarah):

waiit "You have an equally likely chance of chooseing any integer from 1 through 80. Find the probability of the given event. An odd number is chosen

OpenStudy (ammarah):

would it be any odd number?

OpenStudy (tkhunny):

Are there just as many odd numbers as there are even numbers? Done!

OpenStudy (ammarah):

for probability do i write the ratio?

OpenStudy (ammarah):

what does that mean? yes

OpenStudy (ammarah):

i got 5/80 as an answer and i changed it to a decimal

OpenStudy (tkhunny):

p(odd from 1-80) = 40/80 = 1/5 = 0.50 = 50% Odds of odd from 1-80 are 40:40

OpenStudy (ammarah):

how about for a multiple of 3

OpenStudy (ammarah):

im sorry last one so i can make sure im doing this right

OpenStudy (ammarah):

a perfect square

OpenStudy (ammarah):

a factor of 200 is chosen!!